# Two dimensional dimers beyond planarity

### v2

@@ 1 Introduction
$\sum\limits_{x\in\mathbb{S}_{L,K}}\frac{\mathcal{C}_{L,K}(o,x)}{|\mathbb{S}_{L,K}|}\leq c\sqrt{\frac{K}{\log L}}$ (1.4)
Contrary to (1.4) the Cesaro sum of monomer-monomer correlations is known to be uniformly positive in the $d$-dimensional torus if $d\geq 3$ [30]. Note that both our theorems also hold for the planar torus, corresponding to the case $K=1$, and even in this case Theorem 1.1 is new when $\rho>0$, being Kasteleyn’s theorem inapplicable.
Discussion.
@@
When $N=1$ the monomer double dimer model (and the general class of models introduced in Section 3 below) is expected to undergo a Kosterlitz-Thouless phase transition with respect to the variation of the monomer activity on two dimensional graphs. This means that the connection probability exhibits exponential decay for large enough monomer activity (which has been proved in [4]) and polynomial decay for small enough (but possibly positive) monomer activity. We may hope that the spin representation may help us to import techniques from spin systems in the framework of perfect matchings to make further progress. For example, one may hope to apply the recent methods of [1, 12] or the technique of [16] to show that the connection probability decays not faster than polynomially on two dimensional graphs if the monomer activity is small enough and $N=1$, or to make further progress with the dimer problem (with or without monomers) in dimensions three and higher [28, 30].
Paper organization.
@@
In Section 2 we introduce the complex spin representation and state our main theorem for spins. In Section 3 we introduce our generalized double dimer model, discuss some special cases and make a connection between dimers and spins. In Section 4 we introduce the reflection positivity property and discuss its consequences. In Section 5 we present some probabilistic estimates, introduce a central quantity, the magnetisation, and present a perturbative analysis in the limit of vanishing external field which allows us to lower bound the magnetisation by the Cesaro sum of two point functions at zero external field. Section 6 is the central part of our paper, here we introduce our central tool, a weak version of the Cauchy Schwarz for our complex measure. In Section 7 we present the proof of our main theorems and describe the generalisation to arbitrary number of colours. Finally, in Section 8 we present the proof of our proposition which relates spins and dimers.
@@
Unless otherwise specified, the constants $c$, $c_{i}$, $i=1,2,\ldots$, appearing in the text will always be positive, finite and will possibly depend on the parameter $\beta$, which is introduced below and corresponds to the inverse monomer activity in the monomer double dimer model. Their value may change from line to line. Any further dependence of such constants from the other parameters will be specified explicitly. In the whole paper the graph $G=(V,E)$ will always be a undirected finite bipartite graph. Moreover, $d(x,y)$ will be used to denote the graph distance. We use the notation $\mathbb{N}_{0}=\{0,1,\ldots\}$, =$\mathbb{N}=\{1,2,\ldots\}$, $\mathbb{R}^{+}_{0}=\{x\in\mathbb{R}:x\geq 0\}$. Finally, we use the abbreviations ‘LHS’ for ‘left-hand side’ and ‘RHS’ for ‘right-hand side’.
@@ 2 The spin representation
We let $G=(V,E)$ be a undirected finite bipartite graph with $o\in V$ a prescribed vertex. We let $V^{e}$ (resp. $V^{o}$) be the set of vertices having even (resp. odd) graph distance from $o$. We let $U:\mathbb{N}_{0}\mapsto\mathbb{R}^{+}_{0}$ be a weight function, which defines our model. In the whole paper we assume that the non negative sequence $\big{(}U(n)\big{)}_{n\in\mathbb{N}_{0}}$ is sumable, that $U(n)>0$ for some $n>0$, and that $U(n)\leq 1$ for any $n\in\mathbb{N}_{0}$. We introduce a ‘local’ spin space $S=[0,2\pi)^{2}$. We define for each $s=(s^{1},s^{2})\in S$ the function
$g(s):=\frac{1}{(2\pi)^{2}}\sum\limits_{n=0}^{\infty}U(n)e^{-n\,i\,(s^{1}-s^{2})}.$ (2.1)
We denote by $\Omega_{s}:=S^{V}$ the configuration space and define for each $\boldsymbol{s}=(s_{z})_{z\in V}\in\Omega_{s}$, with $s_{z}=(s_{z}^{1},s_{z}^{2})$,
$\boldsymbol{g}(\boldsymbol{s}):=\prod_{i\in V^{e}}g(s_{i})\prod_{j\in V^{o}}\overline{g}(s_{j})$ (2.2)
@@
and interpret $s^{1}$, $s^{2}$ as the angles of the two components of the spin. We define the function
$H(\boldsymbol{s})$ :=$:=\sum\limits_{\{i,j\}\in E}e^{i(s_{i}^{1}-s_{j}^{1})}+\sum\limits_{\{i,j\}\in E}e^{-i(s_{i}^{2}-s_{j}^{2})}+\,h\sum\limits_{k\in V}\,2\,\cos(rs^{1}_{k}).$ (2.3)
In the previous expressions (and in the whole paper) we use the convention that the sum over indirected edges $\{i,j\}$ the first vertex, $i$, is always even and the second vertex, $j$, is always odd. Moreover, $h,r\geq 0$ are some parameters. We define the complex measures assigning to each measurable complex function $f:\Omega_{s}\mapsto\mathbb{C}$ the weight
@@
$\langle f\rangle_{G,\beta,h,r}$ $:=\frac{1}{Z_{G,\beta,h,r}}\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{g}(\boldsymbol{s})\,\,e^{\beta H(\boldsymbol{s})}\,\,f(\boldsymbol{s})$ (2.4)
$Z_{G,\beta,h,r}$ $:=\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{g}(\boldsymbol{s})\,\,e^{\beta H(\boldsymbol{s})},$ (2.5)
where we integrate with respect to a product of Lebeasgue measures in $S^{V}$ and $\beta\geq 0$ is a parameter, the inverse temperature, which will play the role of the inverse monomer activity in the monomer double dimer model. Clearly, if $h=0$ the operator $\langle\,\,\rangle_{G,\beta,h,r}$ does not depend on the value of $r$.
@@
Our spin representation is invariant with respect to global spin rotations. Indeed, due to our choice of (2.1), if we add the same constant to the ‘angles’ $s_{x}^{1}$ and $s_{x}^{2}$ for any $x\in V$ the integrand in (2.4) does not change. This feature will turn out to be essential for the proof of our main theorem, and seems to be the analogous of the invariance under rotation property in Spin O(N) models with $N\geq 2$.
@@
Our ultimate goal is to obtain information on the behaviour of the two-point function when $h=0$. We consider two different two-point functions, corresponding to the cases $\ell=1$ and $\ell=2$ in equation (2.6) below. These are related to two different important quantities in our general double dimer model (we explore this connection in Proposition 3.3 below). We will obtain information on such two point functions by considering the limit $h\rightarrow 0$ of our spin system when $r=1$ and $r=2$ in (2.3) respectively. We now state our main theorem about decay of correlations in the spin system (2.4). When the graph $G$ is the slab torus $\mathbb{S}_{L,K}$, we use the subscript ${}_{L,K}$ rather than ${}_{\mathbb{S}_{L,K}}$ in order to simplify the notation.
@@
Fix $\ell\in\{1,2\}$, let $\beta\geq 0$ be arbitrary. There exists $c=c(\beta)\in(0,\infty)$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq\sqrt{\log(L)}$, we have
$0\leq\sum\limits_{x\in\mathbb{S}_{L,K}}\frac{1}{|\mathbb{S}_{L,K}|}\langle\,\cos(\ell\,s_{o}^{1})\cos(\ell\,s_{x}^{1})\,\rangle_{L,K,\beta,0,0}\leq c\sqrt{\frac{K}{\log L}}$ (2.6)
@@
3 A generalized double dimer model
We introduce a generalisation of the double dimer model, in which each vertex is touched by the same number of dimers for each type ($1$=blue and $2=$red) and this number is an arbitrary non-negative integer. This generalisation includes the monomer double dimer model, in which such a number is either zero or one, and the double dimer model, in which such a number is precisely one. In Section 7.1 below we deal with the case of arbitrary number of colours.

### v3

@@ 1.3.1 Dimer model
$\sum\limits_{x\in\mathbb{S}_{L,K}}\frac{\mathcal{C}_{L,K}(o,x)}{|\mathbb{S}_{L,K}|}\leq c\sqrt{\frac{K}{\log(L)}}$ (1.2)
The behaviour of the monomer-monomer correlation was investigated in [18] with the aid of a general perturbation theory for Pfaffians. In this paper it was conjectured the polynomial decay of this quantity with exponent $\frac{1}{2}$ after taking the thermodynamic limit, namely
$\lim\limits_{L\rightarrow\infty}\mathcal{C}_{L,1}(o,x)\sim|x|^{-\frac{1}{2}}$ (1.3)
in the limit of large $|x|$. This conjecture, known as the Fisher-Stephenson conjecture, was later solved by Dubédat (see [15, Section 8]). See also, for example, [17, 26, 44, 56] for further rigorous and non-rigorous studies of monomer correlations.
The rigorous results of these works strongly rely on Kasteleyn’s theorem, namely on the integrability property of the model on planar graphs. Without such integrability features, it is difficult to expect sharp information on the precise decay of the correlation function. The novelty of our result is that it is possible to obtain non-trivial information on this model in two dimensions—namely, the decay to zero of the monomer–monomer correlation function on average, and thus the absence of long-range order in two dimensions—without relying on integrability properties. Contrary to (1.2) the Cesaro sum of monomer-monomer correlations is known to be uniformly positive in the $d$-dimensional torus if $d\geq 3$ [50].
Note that our theorem states the absence of long-range order still holds for slabs whose thickness may grow with $L$.
1.3.2 Monomer double-dimer model and spatial random permutations
The presence of a positive density of monomers destroys planarity and prevents the use of Kasteleyn’s theorem. As a consequence, the monomer double–dimer model is highly non-trivial and has remained completely unexplored. Our main theorem provides the first rigorous result for this model in two dimensions: it establishes the absence of macroscopic loops, the decay of correlations for any monomer density, and thus the absence of long-range order.
The configuration space of the monomer double-dimer model is the set
$\Omega:=\{\omega=(M,d_{1},d_{2})\,\,:\,\,M\subset V,(d_{1},d_{2})\in\mathcal{D}_{G}(M)\times\mathcal{D}_{G}(M)\}.$
We refer to the first element of the triplet $\omega=(M,d_{1},d_{2})\in\Omega$ as a set of monomers, and to the second and third element as set of dimers (respectively: blue and red dimers). For example, a configuration of the monomer double-dimer model with four monomers in the box of $\mathbb{Z}^{2}$ can be seen on the left of Figure 1.1. As one can see from the figure, any configuration $(M,d_{1},d_{2})\in\Omega$ can be viewed as a collection of disjoint self-avoiding loops, with loops having length at least four and corresponding to an alternation of blue and red dimers or having length two and corresponding to the superposition of one blue and one red dimer on the same edge.
We define a probability measure on $\Omega$ which assigns to each configuration $\omega=(M,d_{1},d_{2})\in\Omega$ the weight
$\forall\omega\in\Omega\qquad\mathbb{P}_{G,\rho}(\omega):=\frac{\rho^{|M|}}{\mathbb{Z}_{G,\rho}},$ (1.4)
where $\rho\geq 0$ is a parameter, the monomer activity, and $|M|$ denotes the cardinality of the set of monomers $M$, and $\mathbb{Z}_{G,\rho}$ is a normalising constant. We use $\mathbb{E}_{G,\rho}$ to denote the expectation with respect to (1.4).
In the special case $\rho=0$ the monomers are suppressed almost surely, hence one recovers the double-dimer model, corresponding to the superposition of two dimer covers of $G$ drawn independently and uniformly at random ([16, 30]).
In order to state our main theorem for the monomer double-dimer model we need to introduce two important observables. The first quantity is the loop length. We define $L_{x}={L}_{x}(\omega)$ the subgraph of $G$ corresponding to the loop of $\omega$ touching the vertex $x\in V$ in the configuration $\omega\in\Omega$ and set $L_{x}(\omega)=\emptyset$ if the configuration $\omega$ has a monomer at $x$ . We denote by $|{L}_{x}|=|{L}_{x}|(\omega)$ the length of the loop $L_{x}(\omega)$, namely the number of vertices which are touched by such a loop, and set $|{L}_{x}|(\omega)=0$ if $\omega$ has a monomer at $x$.
The second interesting quantity is the two-point function, corresponding to ratio of the total weight of configurations with a walk having two prescribed vertices as end-points and the partition function. More formally, if $x$ and $y$ are two vertices having odd distance, then we define $\Omega_{x,y}$ as the set of triplets $\omega=(M,d_{1},d_{2})$ such that $M\subset V\setminus\{x,y\}$, $d_{1}\in\mathcal{D}_{G}(\{x,y\}\cup M)$, and $d_{2}\in\mathcal{D}_{G}(M)$. If, instead, $x$ and $y$ have even distance, we define $\Omega_{x,y}$ as the set of triplets $\omega=(M,d_{1},d_{2})$ such that $M\subset V\setminus\{x,y\}$, $d_{1}\in\mathcal{D}_{G}(\{x\}\cup M)$, and $d_{2}\in\mathcal{D}_{G}(\{y\}\cup M)$. As one can see on the right of Figure 1.1, any configuration $\omega\in\Omega_{x,y}$ can be viewed as a collection of monomers, mutually disjoint self-avoiding loops and a self-avoiding walk with $x$ and $y$ as end-point. We define the two-point function for the monomer double-dimer model
$\mathbb{G}_{G,\rho}(x,y):=\frac{\sum\limits_{\omega=(M,d_{1},d_{2})\in\Omega_{x,y}}\rho^{|M|}}{\mathbb{Z}_{G,\rho}}$
The first claim in our theorem states that the expected loop length divided by the volume converges to zero in the limit of large slabs. The second claim states the absence of macroscopic loops. The third claim states the decay to zero of the two-point function and thus the absence of long-range order in the monomer double-dimer model. Our results hold for any value of the monomer activity.
Theorem 1.2.
Suppose that $\rho\geq 0$. There exists $c=c(\rho)\in(0,\infty)$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq\sqrt{\log(L)}$, and for any $\varepsilon>0$
$\frac{\mathbb{E}_{L,K,\rho}\big(|L_{o}|\big)}{|\mathbb{S}_{L,K}|}$ $\leq c\sqrt{\frac{K}{\log(L)}},$ (1.5)
$\mathbb{P}_{L,K,\rho}\big(|L_{o}|>\varepsilon|\mathbb{S}_{L,K}|\big)$ $\leq c\varepsilon^{-1}\sqrt{\frac{K}{\log(L)}},$ (1.6)
$\sum\limits_{x\in\mathbb{S}_{L,K}}\frac{\mathbb{G}_{L,K,\rho}(o,x)}{|\mathbb{S}_{L,K}|}$ $\leq c\sqrt{\frac{K}{\log(L)}}.$ (1.7)
Since $\mathbb{E}_{L,K,\rho}\big(|L_{o}|\big)=\sum_{x\in\mathbb{S}_{L,K}}\mathbb{P}_{L,K,\rho}(o\leftrightarrow x)$, where $\{x\leftrightarrow y\}\subset\Omega$ is the event that there exists a loop connecting $x$ and $y$, our theorem implies that the probability that two distant vertices are visited by the same loop is not uniformly positive, contrary to what happens in dimension three and higher and small monomer activity.
Note that, similarly to Theorem 1.2, the absence of long-range order and of macroscopic loops holds true for slabs whose size can grow even slightly faster than $O(L^{2})$.
Moreover our results also hold for a generalisation of the model in which an additional multiplicative factor $N^{\#\mbox{ \small loops}}$ is added to the measure, see Section 8.1 below.
Spatial random permutations.
The model of spatial random permutations on a graph $G=(V,E)$ is defined as follows. Let $\Omega^{per}$ be the set of permutations $\pi$ of $V$ such that, for each $x\in V$,
$d_{G}(x,\pi(x))\in\{0,1\},$
where $d_{G}(x,y)$ denotes the graph distance on $G$. We define a probability measure ${P}^{per}_{G,\rho}$ on $\Omega^{per}$ by setting, for each $\pi\in\Omega^{per}$,
$P_{G,\rho}^{per}(\pi)=\frac{\rho^{|F(\pi)|}}{Z^{per}_{G,\rho}},$ (1.8)
where
$F(\pi):=\{x\in V:\pi(x)=x\}$
is the set of fixed points of the permutation $\pi$, $\rho\geq 0$ is a parameter, and $Z^{per}_{G,\rho}$ is the normalising constant.
As illustrated in Figure 1.1, there is a one-to-one correspondence between elements of $\Omega$ (i.e., realisations of the monomer double-dimer model) and elements of $\Omega^{per}$. Moreover, the measure (1.8) assigns to each permutation the same weight as (1.4) assigns to the corresponding monomer double-dimer realisation. Thus, the two models are equivalent reformulations of each other; any result stated in this paper for the monomer double-dimer model can be restated for spatial random permutations on a lattice.
Open problems.
@@ 1.3.2 Monomer double-dimer model and spatial random permutations
The monomer double-dimer model is expected to undergo a Kosterlitz–Thouless phase transition as the monomer activity varies on two-dimensional graphs. This means that the connection probability and the two-point function exhibit exponential decay for large enough monomer activity (which has been proved in [7], where non-trivial estimates of the critical threshold are obtained) and polynomial decay for small enough (but possibly positive) monomer activity. We may hope that the spin representation may help us to import techniques from spin systems in the framework of perfect matchings to make further progress. For example, one may hope to apply the recent methods of [1, 36] or the technique of [19] to show that the connection probability decays not faster than polynomially on two-dimensional graphs if the monomer activity is small enough. Another remarkable open problem is proving the existence of macroscopic loops in dimension three and higher when the monomer activity is strictly positive but sufficiently small, a result which has been proved only at zero monomer activity [47]. The same discussion applies to spatial random permutations, since the model is equivalent to the monomer double-dimer model. For a closely related variant of the spatial random permutations considered here, the occurrence of a Kosterlitz–Thouless phase transition in two dimensions has been conjectured and investigated numerically in [6].
1.4 Paper organisation and notation
@@
In Section 2 we introduce the complex spin representation and state our main theorem for spins. In Section 3 we introduce the multi occupancy double-dimer model, discuss some special cases and make a connection between dimers and spins. In Section 4 we introduce the reflection positivity property and discuss its consequences. In Section 5 we present some probabilistic estimates, introduce a central quantity, the magnetisation, and present a perturbative analysis in the limit of vanishing external field which allows us to lower bound the magnetisation by the Cesaro sum of two point functions at zero external field. Section 6 is the central part of our paper, here we introduce our central tool, a weak version of the Cauchy Schwarz for our complex measure (a Bogoliubov-type inequality). In Section 7 we present the proof of our main theorems. Finally, in Section 8 we describe the generalisation to arbitrary number of colours and we present the proof of our propositions which relates spins and dimers.
@@
Unless otherwise specified, the constants $c$, $c_{i}$, $i=1,2,\ldots$, appearing in the text will always be positive, finite and will possibly depend on the parameter $\beta$, which is introduced below and corresponds to the inverse monomer activity in the monomer double-dimer model. Their value may change from line to line. Any further dependence of such constants from the other parameters will be specified explicitly. In the whole paper the graph $G=(V,E)$ will always be an undirected finite bipartite graph. We write $x\sim y$ whenever $\{x,y\}\in E$. Moreover, $d(x,y)$ will be used to denote the graph distance. We use the notation $\mathbb{N}_{0}=\{0,1,\ldots\}$, $\mathbb{N}=\{1,2,\ldots\}$, $\mathbb{R}^{+}_{0}=\{x\in\mathbb{R}:x\geq 0\}$. Finally, we use the abbreviations ‘LHS’ for ‘left-hand side’ and ‘RHS’ for ‘right-hand side’.
@@ 2 The spin representation
We let $G=(V,E)$ be an undirected finite bipartite graph with $o\in V$ a prescribed vertex. We let $V^{e}$ (resp. $V^{o}$) be the set of vertices having even (resp. odd) graph distance from $o$. We let $U:\mathbb{N}_{0}\mapsto\mathbb{R}^{+}_{0}$ be a weight function, which defines our model. In the whole paper we assume that the non negative sequence $\big(U(n)\big)_{n\in\mathbb{N}_{0}}$ is summable, that $U(n)>0$ for some $n>0$, and that $U(n)\leq 1$ for any $n\in\mathbb{N}_{0}$. We introduce a ‘local’ spin space $\Xi=[0,2\pi)^{2}$. We denote by $\Omega_{s}:=\Xi^{V}$ the configuration space and define for each $\boldsymbol{s}=(s_{z})_{z\in V}\in\Omega_{s}$, with $s_{z}=(s_{z}^{1},s_{z}^{2})$ the spin at $z$, namely the vector function
$S_{z}=(S_{z}^{1},S_{z}^{2}),$
where the function $S_{z}^{k}:\Omega_{s}\mapsto\mathbb{C}$ is defined for each $k\in\{1,2\}$ as
[…]
@@
and interpret $s^{1}_{z}$, $s^{2}_{z}$ as the angles of the two components of the spin $S_{z}$. We define for each $\boldsymbol{s}\in\Omega_{s}$ the function
$\gamma_{z}(\boldsymbol{s}):=\frac{1}{(2\pi)^{2}}\sum\limits_{n=0}^{\infty}U(n)\big(\overline{S_{z}^{1}(\boldsymbol{s})S_{z}^{2}(\boldsymbol{s})}\big)^{n},$ (2.2)
and
$\boldsymbol{\gamma}(\boldsymbol{s}):=\prod_{z\in V}\gamma_{z}(\boldsymbol{s}).$ (2.3)
We define the Hamiltonian function
$H(\boldsymbol{s})$ $:=\sum\limits_{\{i,j\}\in E}\big(S_{i}^{1}(\boldsymbol{s})S_{j}^{1}(\boldsymbol{s})+S_{i}^{2}(\boldsymbol{s})S_{j}^{2}(\boldsymbol{s})\big)+\,h\sum\limits_{k\in V}\,2\,\cos(rs^{1}_{k}),$ (2.4)
where $h,r\geq 0$ are some parameters. When using the spin representation of dimer models we will later restrict to $r\in\{1,2\}$, but the definition of the spin measure makes sense for any $r\geq 0$.
We define the complex measures that assign to each measurable complex function $f:\Omega_{s}\to\mathbb{C}$ the weight
@@
$\langle f\rangle_{G,\beta,h,r}$ $:=\frac{1}{Z^{spin}_{G,\beta,h,r}}\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{\gamma}(\boldsymbol{s})\,e^{\beta H(\boldsymbol{s})}\,f(\boldsymbol{s}),$ (2.5)
$Z^{spin}_{G,\beta,h,r}$ $:=\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{\gamma}(\boldsymbol{s})\,e^{\beta H(\boldsymbol{s})},$ (2.6)
where the integration is taken with respect to the product of Lebesgue measures on $\Omega_{s}$. Here $\beta\geq 0$ denotes the inverse temperature, which in the monomer double-dimer model plays the role of the inverse monomer activity. Using the correspondence between the spin models and graphical configurations (see Proposition 3.11), we see that $Z^{spin}_{G,\beta,h,r}$ is real and strictly positive as long as $G$ admits a dimer configuration, namely $\mathcal{D}_{G}$ is non-empty. The parameter $h$ plays the role of the external field intensity for our spin measure. The parameter $r$, on the other hand, specifies the type of external field. Its value determines which type of two-point function can be extracted in the case $h=0$, by taking the limit $h\to 0$ in an appropriate way. For instance, when $r=2$ we obtain information about the probability that a loop connects two distant vertices, whereas when $r=1$ we recover the classical two-point function, namely the ratio between the partition function with a walk connecting two points and the partition function restricted to loop configurations only.
Remark 2.1.
@@
When $h=0$, the spin representation is invariant under global spin rotations. Indeed, by our definition (2.1), adding the same constant to the ‘angles’ $s_{x}^{1}$ and $s_{x}^{2}$ for every $x\in V$ leaves the integrand in (2.5) unchanged. This invariance will play a crucial role in the proof of our main theorem, and it can be viewed as the analogue of the rotation invariance in $O(N)$ spin models with $N\geq 2$.
Moreover, due to our definition (2.1), (2.3) satisfies a complex chessboard property, namely that odd sites are associated with functions corresponding to the complex conjugate of those associated with even sites. This feature is essential for establishing reflection positivity.
@@
Our ultimate goal is to obtain information on the behaviour of the two-point function when $h=0$. We focus on two distinct two-point functions, corresponding to the cases $r=1$ and $r=2$ in equation (2.7) below. These functions are related to two important quantities in our general double-dimer model (see Proposition 3.15 for details).
To extract information on these two-point functions, we consider the limit $h\to 0$ of our spin system for $r=1$ and $r=2$ in (2.4), respectively. In (2.4) the parameter $r$ controls the ‘type’ of the external field. We are now in position to state our main theorem on the decay of correlations in the spin system (2.5).
@@
Fix $r\in\{1,2\}$, let $\beta\geq 0$ be arbitrary. There exists $c=c(\beta)\in(0,\infty)$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq\sqrt{\log(L)}$, we have
$0\leq\sum\limits_{x\in\mathbb{S}_{L,K}}\frac{1}{|\mathbb{S}_{L,K}|}\langle\,\cos(r\,s_{o}^{1})\cos(r\,s_{x}^{1})\,\rangle_{L,K,\beta,0,0}\leq c\sqrt{\frac{K}{\log(L)}}$ (2.7)
@@
3 Multi-occupancy double-dimer model
We introduce a generalisation of the monomer double-dimer model in which each vertex is incident to the same number of dimers of each type (1=1=blue, 2=2=red), where this number is an arbitrary non-negative integer. By introducing suitable pairing variables that match, at each vertex, every blue dimer with a corresponding red dimer, one can still interpret configurations in this more general setting as collections of loops.
This framework encompasses both the monomer double-dimer model, where the number is restricted to 0 or $1$, and the standard double-dimer model, where it is exactly $1$.
In Section 8.1 we further extend this construction to the case of an arbitrary (even) number of colours.